### v1

# TreeFive.bib

Abstract.
We prove that a tree with a vertex of degree at least five must be missing a connected partition of some type and therefore its chromatic symmetric function cannot be $e$-positive. We prove that this also holds for a tree with a vertex of degree four as long as it is not adjacent to any leaf. This brings us very close to the conjecture by Dahlberg, She, and van Willigenburg of non-$e$-positivity for all trees with a vertex of degree at least four. We also prove that spiders with four legs cannot have an $e$-positive chromatic symmetric function.
Key words and phrases:
chromatic symmetric function, connected partition, cut vertex, e-positive, proper colouring, spider graph, Stanley–Stembridge conjecture, tree isomorphism conjecture
2020 Mathematics Subject Classification:
Primary 05E05; Secondary 05C05, 05C15
Department of Mathematics, MIT
1. Introduction
When Stanley [chromsym] defined the chromatic symmetric function $X_{G}$ of a graph $G$, he posed two monumental problems that have garnered tremendous interest. The first is the Stanley–Stembridge conjecture [stanstem], which asserts that $X_{G}$ is $e$-positive whenever $G$ is the incomparability graph of a $(\bm{3}+\bm{1})$-free poset. Many authors proved $e$-positivity for particular subclasses of such graphs [csfvertex, chrombounce3, chromstoe, chromqrook, trilad, chromnsym, qforesttriples] and studied connections to Hessenberg varieties [chromhesssplitting, dothessenberg, hessenberghopf, stanstemhess, chromquasihessenberg] and Hecke algebras [chromhecke, chromcharacters]. The second is the tree isomorphism conjecture, which asserts that trees $T_{1}$ and $T_{2}$ can have $X_{T_{1}}=X_{T_{2}}$ only if they are isomorphic. This is intriguing because all trees with $n$ vertices have the same chromatic polynomial. Many authors proved that the chromatic symmetric function determines certain invariants and therefore distinguishes particular subclasses of trees [propcatchrom, chromtreepolyinv, markedgraphs, propqcatchrom, algochromtrees, catchromiso, distinguishingtrees, chromequal].
Dahlberg, She, and van Willigenburg [chromspostrees] considered a hybrid of these two questions by asking which trees $T$ have that $X_{T}$ is $e$-positive. Many authors [eposclawcon, spider4m22m1, spiderbroom, spiderepos] proved $e$-positivity or non-$e$-positivity for particular subclasses of trees. Dahlberg, She, and van Willigenburg checked the following conjecture for all trees with at most $12$ vertices.
Conjecture 1.1.
[chromspostrees, Conjecture 6.1] If $T$ is a tree with a vertex of degree $d\geq 4$, then $X_{T}$ is not $e$-positive.
A powerful tool for studying $e$-positivity is the following necessary condition by Wolfgang.
Proposition 1.2.
[conparts, Proposition 1.3.3] A connected graph $G$ must have a connected partition of every type in order for $X_{G}$ to be $e$-positive.
Note that this condition is not sufficient because for example, for every $m\geq 1$ the spider graph $G=S(2m+2,1,1)$ has a connected partition of every type but $X_{G}$ is not $e$-positive. Dahlberg, She, and van Willigenburg proved the following result.
Theorem 1.3.
[chromspostrees, Theorem 4.1] If $G$ is an $n$-vertex connected graph with a cut vertex whose deletion produces a graph with $d\geq\log_{2}n+1$ connected components, then $G$ is missing a connected partition of some type and therefore $X_{G}$ is not $e$-positive.
In particular, if $T$ is an $n$-vertex tree with a vertex of degree $d\geq\log_{2}n+1$, then $X_{T}$ is not $e$-positive. Zheng strengthened this result considerably.
Theorem 1.4.
[spiderepos, Theorem 3.13] If $G$ is a connected graph with a cut vertex whose deletion produces a graph with at least $6$ connected components, then $G$ is missing a connected partition of some type.
Our main result lowers this bound further.
Theorem 1.5.
Let $G$ be a connected graph with a cut vertex whose deletion produces a graph with either at least $5$ connected components, or $4$ connected components that each have at least two vertices. Then $G$ is missing a connected partition of some type.
In particular, if $T$ is a tree with a vertex of degree at least $5$, or with a vertex of degree $4$ that is not adjacent to any leaf, then $T$ is missing a connected partition of some type and therefore $X_{T}$ is not $e$-positive.
Note that there exists a tree $T$ with a vertex of degree $4$ that is adjacent to a leaf, and which has a connected partition of every type. Therefore, in this sense, Theorem 1.5 cannot be improved. In fact, we will show in Proposition 3.19 that for every $m\geq 1$ the spider graph $S(6m,6m-2,1,1)$ has a connected partition of every type, so there are infinitely many such examples. This means that to prove non-$e$-positivity of trees with a vertex of degree $4$ in general, we need to know more about the tree to calculate particular coefficients. In the specific case of spiders, we will prove that Conjecture 1.1 holds.
Theorem 1.6.
If $S$ is a spider with four legs, then $X_{S}$ is not $e$-positive.
Note that there exist infinite families of spiders with three legs, such as $S(m+1,m,1)$ for $m\geq 1$, that have $e$-positive chromatic symmetric function [chromspostrees, Theorem 3.2].
In the next section, we define chromatic symmetric functions, connected partitions, and spider graphs. In Section 3, we prove Theorem 1.5 by analyzing partial sums of rearrangements of partitions.
2. Background
A composition is a finite sequence of positive integers $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$, which we call the parts of $\alpha$. We denote by
$\text{set}(\alpha)=\{\alpha_{1},\alpha_{1}+\alpha_{2},\ldots,\alpha_{1}+\cdots+\alpha_{\ell-2},\alpha_{1}+\cdots+\alpha_{\ell-1}\}$
the set of (proper) partial sums of $\alpha$. A partition is a composition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ that also satisfies $\lambda_{1}\geq\cdots\geq\lambda_{\ell}$. We write $\lambda\vdash n$ if $\lambda_{1}+\cdots+\lambda_{\ell}=n$. We say that $\alpha$ is a rearrangement of $\lambda$ if every part occurs the same number of times in $\alpha$ and in $\lambda$.
Example 2.1.
The composition $\alpha=(3,4,2,4)$ has $\text{set}(\alpha)=\{3,7,9\}$ and is a rearrangement of the partition $\lambda=(4,4,3,2)\vdash 13$.
Let $G=(V,E)$ be a graph with $n$ vertices. A proper colouring of $G$ is a function $\kappa:V\to\{1,2,3,\ldots\}$ such that whenever vertices $u,v\in V$ are adjacent, we must have $\kappa(u)\neq\kappa(v)$. The chromatic symmetric function of $G$, defined by Stanley [chromsym, Definition 2.1] is the formal power series in infinitely many commuting variables $(x_{1},x_{2},x_{3},\ldots)$ given by
$X_{G}=\sum_{\kappa\text{ proper colouring }}\prod_{v\in V}x_{\kappa(v)}.$
We can write $X_{G}$ uniquely as a linear combination of elementary symmetric functions $e_{\lambda}$, which are indexed by partitions $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})\vdash n$ and defined by
$e_{\lambda}=e_{\lambda_{1}}\cdots e_{\lambda_{\ell}},\text{ where }e_{k}=\sum_{i_{1}\leq\cdots\leq i_{k}}x_{i_{1}}\cdots x_{i_{k}}.$
We say that $X_{G}$ is $e$-positive if all of the coefficients $c_{\lambda}$ in the expansion $X_{G}=\sum_{\lambda\vdash n}c_{\lambda}e_{\lambda}$ are nonnegative. It is also common to say that $G$ itself is $e$-positive. A connected partition of $G$ of type $\lambda\vdash n$ is a collection $\mathcal{S}=\{S_{1},\ldots,S_{\ell}\}$ of pairwise disjoint subsets of $V$ such that
  • $\circ$
    $\bigcup_{i=1}^{\ell}S_{i}=V$,
  • $\circ$
    the induced subgraph $G[S_{i}]$ is connected for every $1\leq i\leq\ell$, and
  • $\circ$
    $|S_{i}|=\lambda_{i}$ for every $1\leq i\leq\ell$.
Proposition 2.2.
[conparts, Proposition 1.3.3] Let $G$ be a connected $e$-positive graph with $n$ vertices. Then for every partition $\lambda\vdash n$, there exists a connected partition of $G$ of type $\lambda$.
A tree is a connected acyclic graph. A leaf of a tree is a vertex of degree one. A cut vertex of a connected graph $G$ is a vertex whose deletion results in a disconnected graph. For a partition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$, we denote by $S(\lambda)$ the graph consisting of paths of lengths $\lambda_{1},\ldots,\lambda_{\ell}$, called legs, joined at a common vertex. Graphs of the form $S(\lambda)$ are called spiders.
Example 2.3.
Figure 1 shows some spiders and their chromatic symmetric functions. The spider $S(1,1,1)$ is missing a connected partition of type $(2,2)$ and is indeed not $e$-positive. The spider $S(3,2,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 7$ and is $e$-positive. The spider $S(4,1,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 7$ but is not $e$-positive, showing that Wolfgang’s necessary condition for $e$-positivity is not sufficient.
Figure 1. Some spiders and their chromatic symmetric functions
$S(1,1,1)$         $S(3,2,1)$         $S(4,1,1)$
$X_{S(1,1,1)}$ $=e_{211}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-2e_{22}}+5e_{31}+4e_{4}$
$X_{S(3,2,1)}$ $=e_{2221}+2e_{3211}+5e_{322}+4e_{331}+12e_{421}+5e_{43}+4e_{511}+13e_{52}+11e_{61}+7e_{7}$
$X_{S(4,1,1)}$ $=e_{2221}+4e_{3211}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-3e_{322}}+10e_{331}+10e_{421}+17e_{43}+4e_{511}+3e_{52}+11e_{61}+7e_{7}$
3. Proof of main result
We first fix notation that we will use throughout this section. Let $G$ be an $n$-vertex connected graph with a cut vertex $v$ whose deletion produces a graph with at least $3$ connected components, and denote their sizes
$a\geq b\geq c_{1}\geq\cdots\geq c_{k}\geq 1,\text{ where }k\geq 1.$
See Figure 2. Let $c=c_{1}+\cdots+c_{k}$. We will always assume that $c\geq 2$. Note that
$n=a+b+c+1\geq 2b+c+1.$
We now state our main result using this notation.
Theorem 3.1.
Suppose that $c\geq 2$. If any of the following hold, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
  1. (1)
    We have $b\leq 2c-2$.
  2. (2)
    We have $c\geq c_{1}+1$ and $b=2c-1$.
  3. (3)
    We have $2c\leq b\leq\frac{c^{2}}{2}$.
  4. (4)
    We have $c\geq c_{1}+2$ and $b\geq\frac{c^{2}}{2}$.
In particular, if $c\geq c_{1}+2$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Remark 3.2.
The condition that $c\geq c_{1}+2$ is precisely that the deletion of $v$ produces either at least $5$ connected components, or $4$ connected components that each have at least two vertices, so Theorem 3.1 implies Theorem 1.5. We cannot remove this hypothesis because the spider graph $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$.
Remark 3.3.
The condition that $c\geq c_{1}+1$ is precisely that the deletion of $v$ produces at least $4$ connected components. We cannot remove this hypothesis because the spider graph $S(5,3,2)$ has a connected partition of every type $\lambda\vdash 11$ and is in fact $e$-positive.
Remark 3.4.
We cannot remove the hypothesis that $c\geq 2$ because the family of spider graphs $S(m+1,m,1)$ all have a connected partition of every type $\lambda\vdash 2m+3$ and are in fact $e$-positive.
We begin by describing another way to think about connected partitions.
Lemma 3.5.
Let $\lambda\vdash n$ be a partition with all parts at least $(c_{1}+1)$. If every rearrangement $\alpha$ of $\lambda$ has a partial sum $s\in\text{set}(\alpha)$ that lies in the interval $I=\{b+1,b+2,\ldots,b+c-1,b+c\}$, then $G$ has no connected partition of type $\lambda$.
Proof.
Suppose that $G$ has a connected partition $\mathcal{S}$ of type $\lambda$. We will find a rearrangement $\alpha$ of $\lambda$ that has no partial sum in $I$. Let $A,B,C_{1},\ldots,C_{k}$ be the connected components of $G\setminus v$ whose sizes are respectively $a\geq b\geq c_{1}\geq\cdots\geq c_{k}$ and let $S\in\mathcal{S}$ be the subset with $v\in S$. If any $C_{i}\nsubseteq S$, then letting $u\in C_{i}\setminus S$, the subset $S^{\prime}\in\mathcal{S}$ with $u\in S^{\prime}$ would have $|S^{\prime}|\leq c_{i}\leq c_{1}$, which is impossible because every part of $\lambda$ is at least $(c_{1}+1)$. So every $C_{i}\subseteq S$. This means that $\mathcal{S}$ is of the form $\mathcal{S}=\{A_{1},\ldots,A_{i},S,B_{1},\ldots,B_{j}\}$ where $A_{i^{\prime}}\subseteq A$ for $1\leq i^{\prime}\leq i$ and $B_{j^{\prime}}\subseteq B$ for $1\leq j^{\prime}\leq j$. We now take $\alpha$ to be the composition
$\alpha=(|B_{1}|,\ldots,|B_{j}|,|S|,|A_{1}|,\ldots,|A_{i}|).$
By definition, $\alpha$ is a rearrangement of $\lambda$, and we have
$\alpha_{1}+\cdots+\alpha_{j}$ $=|B_{1}|+\cdots+|B_{j}|\leq|B|=b\text{ and }$
$\alpha_{1}+\cdots+\alpha_{j+1}$ $=|B_{1}|+\cdots+|B_{j}|+|S|\geq|B|+|C_{1}|+\cdots+|C_{k}|+1=b+c+1.$
Therefore, no partial sum of $\alpha$ lies in the interval $I=\{b+1,\ldots,b+c\}$. ∎
Figure 2. A graph $G$ with a cut vertex $v$ joining at least $3$ connected components
[…]
We can now prove Part 4 of Theorem 3.1, the case where $b\geq\frac{c^{2}}{2}$.
Proposition 3.6.
If $c\geq c_{1}+2$ and $b\geq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
The idea is to take a partition $\lambda\vdash n$ with all parts equal to $(c-1)$ or $c$. The well-known Frobenius coin problem tells us that every integer $n\geq(c-1)(c-2)$ can be written in the form $n=a_{1}c+a_{2}(c-1)$ for some integers $a_{1},a_{2}\geq 0$. Because $b\geq\frac{c^{2}}{2}$, we have
$n\geq 2b+c+1\geq c^{2}+c+1\geq(c-1)(c-2),$
so we can write $n=a_{1}c+a_{2}(c-1)$ and take $\lambda$ to be $a_{1}$ $c$’s followed by $a_{2}$ $(c-1)$’s. All parts of $\lambda$ are at least $c-1\geq c_{1}+1$. Let $\alpha$ be a rearrangement of $\lambda$. Because $c\geq c_{1}+2\geq 3$, we have $\alpha_{1}\leq c\leq\frac{c^{2}}{2}\leq b$, so there exists some maximal $i$ with $\alpha_{1}+\cdots+\alpha_{i}\leq b$, and then
$b+1\leq\alpha_{1}+\cdots+\alpha_{i}+\alpha_{i+1}\leq b+\alpha_{i+1}\leq b+c.$
Therefore the result follows from Lemma 3.5. ∎
We now describe another way to ensure that all rearrangements have a partial sum in $I$.
Lemma 3.7.
Let $q$ be a positive integer and let $J=\{x,x+1,\ldots,y-1,y\}$, where
(3.1) $x=\left\lceil\frac{b+1}{q}\right\rceil\text{ and }y=\left\lfloor\frac{b+c}{q}\right\rfloor.$
Suppose that $x\geq c_{1}+1$. If $\lambda\vdash n$ is a partition with all parts in the interval $J$, then $G$ has no connected partition of type $\lambda$.
Proof.
All parts of $\lambda$ are at least $x\geq c_{1}+1$ and every rearrangement $\alpha$ of $\lambda$ has
$b+1\leq qx\leq\alpha_{1}+\cdots+\alpha_{q}\leq qy\leq b+c,$
so the result follows from Lemma 3.5. ∎
We now generalize the result of the Frobenius coin problem by seeing which numbers can be written as a sum of numbers in an interval $J$.
Lemma 3.8.
Let $x$ and $y$ be positive integers with $x<y$. Then there exists a partition $\lambda\vdash n$ with all parts in the interval $J=\{x,\ldots,y\}$ as long as
(3.2) $n\geq\left\lceil\frac{x-1}{y-x}\right\rceil x.$
Proof.
For a positive integer $t$, let $tJ=\{tx,tx+1,\ldots,ty-1,ty\}$. First note that every integer $n\in tJ$ has a partition $\lambda\vdash n$ consisting of $t$ integers in $J$. Indeed, for $n=ty$ we can take $\lambda=(y,\ldots,y)$, and for $tx\leq n\leq ty-1$ we can write $n=tq+r$ for some integers $x\leq q\leq y-1$ and $10\leq r\leq t-$ 1, and take $\lambda$ to be $r$ $(q+1)$’s followed by $(t-r)$ $q$’s.
We now see that as $t$ increases, the intervals $tJ$ become closer to each other. Specifically, for $t\geq\left\lceil\frac{x-1}{y-x}\right\rceil\geq\frac{x-1}{y-x}$, we have $ty+1\geq(t+1)x$, which means that as soon as we exit the interval $tJ$, we are already in the interval $(t+1)J$. This means that every integer $n$ satisfying (3.2) is in some interval $tJ$ and has a partition $\lambda\vdash n$ with all parts in $J$. ∎
Example 3.9.
Let $x=7$ and $y=9$, so that $J=\{7,8,9\}$ and we can write the numbers
$\{\underbrace{7,8,9}_{J}\}\cup\{\underbrace{14,15,16,17,18}_{2J}\}\cup\{\underbrace{21,22,23,24,25,26,27}_{3J}\}\cup\{\underbrace{28,29,30,31,32,33,34,35,36}_{4J}\}\cup\cdots$
as a sum of numbers in $J$. We see that once $t\geq\left\lceil\frac{x-1}{y-x}\right\rceil=3$, we no longer have a gap between the intervals $tJ$ and $(t+1)J$, so we are able to express any integer $n\geq 3x=21$.
Proposition 3.10.
If $n\geq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
By Lemma 3.8, there exists a partition $\lambda\vdash n$ with all parts in the interval $J=I=\{b+1,\ldots,b+c\}$. All parts of $\lambda$ are at least $b+1\geq c_{1}+1$ and trivially every rearrangement $\alpha$ of $\lambda$ has the partial sum $\alpha_{1}\in\{b+1,\ldots,b+c\}$, so the result follows from Lemma 3.5. ∎
We can now prove Parts 1 and 2 of Theorem 3.1.
Proposition 3.11.
If $c\geq 2$ and $b\leq 2c-2$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
If $b\leq 2c-2$, then we have
$\left\lceil\frac{b}{c-1}\right\rceil(b+1)\leq 2(b+1)\leq 2b+c+1\leq n,$
so the result follows from Proposition 3.10. ∎
Proposition 3.12.
If $c\geq c_{1}+1$ and $b=2c-1$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
First note that if $c=2$, then $b=3$, $I=\{4,5\}$, and taking $\lambda=(2,\ldots,2)$ if $n$ is even or $\lambda=(3,2,\ldots,2)$ if $n$ is odd, every rearrangement $\alpha$ has $\alpha_{1}+\alpha_{2}=4$ or $5$ and we are done by Lemma 3.5. Now suppose that $c\geq 3$. Let $x=\left\lceil\frac{b+1}{2}\right\rceil=c\geq c_{1}+1$ and $y=\left\lfloor\frac{b+c}{2}\right\rfloor=\left\lfloor\frac{3c-1}{2}\right\rfloor$. By Lemma 3.8, we can find a partition $\lambda\vdash n$ with all parts in $J=\{x,\ldots,y\}$ because
$\left\lceil\frac{x-1}{y-x}\right\rceil x\leq\left\lceil\frac{c-1}{\frac{c-2}{2}}\right\rceil c=\left\lceil 2+\frac{2}{c-2}\right\rceil c\leq 4c\leq 2b+c+1\leq n.$
By Lemma 3.7 with $q=2$, $G$ has no connected partition of type $\lambda$. ∎
Our last case to prove is Part 3, where $2c\leq b\leq\frac{c^{2}}{2}$. We first check several cases by computer.
Proposition 3.13.
If $2\leq c\leq 40$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
By Proposition 3.10, we may assume that $2b+c+1\leq n\leq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, so there are finitely many triples $(b,c,n)$ to consider. For each one, we checked by computer that there exists a positive integer $q$ and a partition $\lambda\vdash n$ with all parts in the corresponding interval $J=\{x,\ldots,y\}$, where $x\geq c+1\geq c_{1}+1$, so the result follows from Lemma 3.7. ∎
Lemma 3.14.
Suppose that we can find a positive integer $q$ such that the integers $x$ and $y$ as in (3.1) satisfy $x\geq c+1$, $x<y$, and
(3.3) $\left\lceil\frac{x-1}{y-x}\right\rceil x\leq 2b+c+1.$
Then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
This follows directly from Lemma 3.7 and Lemma 3.8. ∎
Proposition 3.15.
If $41\leq c\leq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
There are finitely many pairs $(b,c)$ to consider. For each one, we checked by computer that there exists a positive integer $q$ such that $x\geq c+1$, $x<y$, and (3.3) holds. ∎
We now come to the most technical part of the proof, where given any $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, we will find a suitable integer $q$. We first make a remark about this inequality.
Remark 3.16.
If we ignore the floors, ceilings, and ones, we have a rough approximation
$\left\lceil\frac{x-1}{y-x}\right\rceil x\approx\frac{b/q}{c/q}\frac{b}{q}=\frac{b^{2}}{cq}.$
To minimize this, we want $q$ to be large, and to have $x\geq c+1$, we could take $q\approx\frac{b}{c}$, which would give us $\left\lceil\frac{x-1}{y-x}\right\rceil x\approx b$, so this is promising. However, the floors and ceilings mean that the denominator $(y-x)$ may be closer to $(\frac{c}{q}-2)$, and this $(-2)$ term can really hurt us if $q$ is close to $\frac{c}{2}$. Therefore, when $b$ is close to $\frac{c^{2}}{2}$, we will need to make a clever choice of $q$.
Lemma 3.17.
Fix positive integers $b$ and $c$ with $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$. There exists a positive integer $q$ such that, letting $x$ and $y$ be as in (3.1), we have $x\geq c+1$, $x<y$, and the inequality (3.3) holds.
Proof.
We first note that
$y-x=\left\lfloor\frac{b+c}{q}\right\rfloor-\left\lceil\frac{b+1}{q}\right\rceil\geq\left(\frac{b+c}{q}-\frac{q-1}{q}\right)-\left(\frac{b+1}{q}+\frac{q-1}{q}\right)=\frac{c}{q}-2+\frac{1}{q}\geq\frac{c}{q}-2$
because the floors and ceilings round by at most $\frac{q-1}{q}$. We also have
$\left\lceil\frac{x-1}{y-x}\right\rceil x=\left\lceil\frac{\left\lceil\frac{b+1}{q}\right\rceil-1}{y-x}\right\rceil\left\lceil\frac{b+1}{q}\right\rceil\leq\left(\frac{b/q}{y-x}+1\right)\left(\frac{b}{q}+1\right).$
Now we will have a few different cases, depending on the value of $b$.
Case 1.
$2c\leq b\leq\frac{c^{2}}{20}$. We let $q=\left\lfloor\frac{b}{c}\right\rfloor$, which means that $\frac{b}{c}-1\leq q\leq\frac{b}{c}$. Then
$x=\left\lceil\frac{b+1}{q}\right\rceil\geq\left\lceil\frac{b+1}{b}c\right\rceil\geq c+1.$
Because $b\geq 2c$, we have $q\geq 2$ and therefore
$q\geq\frac{2}{3}(q+1)\geq\frac{2}{3}\frac{b}{c}.$
Because $b\leq\frac{c^{2}}{20}$, we have $q\leq\frac{b}{c}\leq\frac{c}{20}$, so $y-x\geq\frac{c}{q}-2\geq 18$ and
$y-x\geq\frac{18}{20}(y-x+2)\geq\frac{18}{20}\frac{c}{q}.$
Now using that $c\geq 500$, we have
$\left\lceil\frac{x-1}{y-x}\right\rceil x$ $\leq\left(\frac{b/q}{y-x}+1\right)\left(\frac{b}{q}+1\right)\leq\left(\frac{20}{18}\frac{b}{c}+1\right)\left(\frac{3}{2}c+1\right)\leq(1.12\frac{b}{c}+1)(1.51c)$
$\leq 1.7b+0.51c+c\leq 1.7b+0.255b+c=1.955b+c\leq 2b+c+1.$
Case 2.
$\frac{c^{2}}{20}\leq b\leq\frac{c^{2}}{4.2}$. We let $q=\left\lfloor\frac{b}{c}\right\rfloor$ again, but our estimates will look a bit different. We still have $x\geq c+1$ as before. Because $b\geq\frac{c^{2}}{20}\geq 25c$, we have $q\geq 25$ and therefore
$q=\frac{25}{26}(q+1)\geq\frac{25}{26}\frac{b}{c}.$
Because $b\leq\frac{c^{2}}{4.2}$, we have $q\leq\frac{b}{c}\leq\frac{c}{4.2}$. Now since $y-x$ is an integer and $y-x\geq\frac{c}{q}-2\geq 2.2$, we must have $y-x\geq 3$ and
$y-x\geq\frac{3}{5}(y-x+2)\geq\frac{3}{5}\frac{c}{q}.$
Now using that $c\geq 500$, we have
$\left\lceil\frac{x-1}{y-x}\right\rceil x$ $\leq\left(\frac{b/q}{y-x}+1\right)\left(\frac{b}{q}+1\right)\leq\left(\frac{5}{3}\frac{b}{c}+1\right)\left(\frac{26}{25}c+1\right)\leq(1.67\frac{b}{c}+1)(1.05c)$
$\leq 1.76b+0.05c+c\leq 1.76b+0.002b+c=1.762b+c\leq 2b+c+1.$
Case 3.
$\frac{c^{2}}{4.2}\leq b\leq\frac{c^{2}}{3.4}$. We let $q=\lfloor 0.46\sqrt{b}\rfloor$. Note that $0.487\leq\frac{1}{\sqrt{4.2}}\leq\frac{\sqrt{b}}{c}\leq\frac{1}{\sqrt{3.4}}\leq 0.543$. This means that
$x=\left\lceil\frac{b+1}{q}\right\rceil\geq\frac{b}{q}\geq\frac{b}{0.46\sqrt{b}}>\frac{\sqrt{b}}{0.487}\geq c,$
so $x\geq c+1$ because $x$ is an integer. We also have
$c\geq\sqrt{3.4}\sqrt{b}\geq\frac{\sqrt{3.4}}{0.46}q\geq 4.008q,$
so because $y-x$ is an integer and $y-x\geq\frac{c}{q}-2>2.008$, we must have $y-x\geq 3$ and
$y-x\geq\frac{3}{5}(y-x+2)\geq\frac{3}{5}\frac{c}{q}.$
Now using that $\sqrt{b}\geq 0.487c\geq 0.487\cdot 500\geq 240$, we have $q\geq 0.46\sqrt{b}-1\geq 0.4558\sqrt{b}$ and
$\left\lceil\frac{x-1}{y-x}\right\rceil x$ $\leq\left(\frac{5}{3}\frac{b}{c}+1\right)\left(\frac{b}{q}+1\right)\leq\left(\frac{5}{3\sqrt{3.4}}\sqrt{b}+1\right)\left(\frac{\sqrt{b}}{0.4558}+1\right)$
$\leq(0.904\sqrt{b}+1)(2.194\sqrt{b}+1)\leq(0.9082\sqrt{b})(2.1982\sqrt{b})\leq 1.997b\leq 2b+c+1.$
Case 4.
$\frac{c^{2}}{3.4}\leq b\leq\frac{c^{2}}{2}$. Let $q_{0}=\left\lfloor\frac{\sqrt{b}}{\sqrt{3.5}}\right\rfloor$. We will either take $q=q_{0}$ or $q=q_{0}-1$. We first show that for one of these values of $q$, we will have $y-x\geq 2$. Note that $\frac{1}{\sqrt{3.4}}\leq\frac{\sqrt{b}}{c}\leq\frac{1}{\sqrt{2}}$, so
$c\geq\sqrt{2}\sqrt{b}\geq\sqrt{2}\sqrt{3.5}q_{0}\geq 2.64q_{0}.$
Also note that because $c\geq 500$, we have $\sqrt{b}\geq\frac{500}{\sqrt{3.4}}\geq 270$ and $q_{0}\geq\frac{270}{\sqrt{3.5}}-1\geq 140$. Let us first consider $q=q_{0}$. Use integer division to write $b+1=c_{0}q_{0}+r_{0}$ for some integers $c_{0}$ and $10\leq r_{0}\leq q_{0}-$ 1. If
$r_{0}\geq 0.38q_{0}\geq 0.37q_{0}+1,$
this will mean that $c+r_{0}-1\geq 2.64q_{0}+0.37q_{0}=3.01q_{0}>3q_{0}$. Therefore,
$b+c=b+1+c-1=c_{0}q_{0}+(c+r_{0}-1)>(c_{0}+3)q_{0},$
so at least three multiples of $q_{0}$, namely $(c_{0}+1)q_{0}$, $(c_{0}+2)q_{0}$, and $(c_{0}+3)q_{0}$, lie strictly between $b+1$ and $b+c$. This means that $x\leq c_{0}+1$ and $y\geq c_{0}+3$, so $y-x\geq 2$.
Now if $r_{0}<0.38q_{0}$, we will consider $q=q_{0}-1$. We have $\sqrt{3.5}q_{0}\leq\sqrt{b}\leq\sqrt{3.5}(q_{0}+1)$, so
$3.53.5q_{0}^{2}\leq b\leq 3.5q_{0}^{2}+7q_{0}+$ 3.5 .
Now the integer $c_{0}$ satisfies $b+1-q_{0}\leq c_{0}q_{0}\leq b+1$, so using that $q_{0}\geq 140$, we have
$3.49q_{0}^{2}\leq 3.5q_{0}^{2}-q_{0}\leq b+1-q_{0}\leq c_{0}q_{0}\leq b+1\leq 3.5q_{0}^{2}+7q_{0}+4.5\leq 3.56q_{0}^{2}.$
This means that in the integer division of $c_{0}$ by $q_{0}$, we have
$c_{0}=3q_{0}+r_{1},\text{ where }0.49q_{0}\leq r_{1}\leq 0.56q_{0}.$
This allows us to calculate the integer division of $b+1$ by $q=q_{0}-1$ as
$b+1=c_{0}q_{0}+r_{0}=c_{0}(q_{0}-1)+3q_{0}+r_{0}+r_{1}=(c_{0}+3)(q_{0}-1)+(r_{0}+r_{1}+3),$
where the remainder $r=r_{0}+r_{1}+3$ satisfies
$0.49(q_{0}-1)\leq r_{1}\leq r\leq 0.38q_{0}+0.56q_{0}+3=0.94q_{0}+3\leq 0.97(q_{0}-1).$
This means that $c+r-1\geq 2.64q_{0}+0.49(q_{0}-1)-1>3(q_{0}-1)$. Therefore,
$b+c=b+1+c-1=(c_{0}+3)(q_{0}-1)+c+r-1>(c_{0}+6)(q_{0}-1),$
so at least three multiples of $(q_{0}-1)$, namely $(c_{0}+4)(q_{0}-1)$, $(c_{0}+5)(q_{0}-1)$, and $(c_{0}+6)(q_{0}-1)$, lie strictly between $(b+1)$ and $(b+c)$. This means that $x\leq c_{0}+4$ and $y\geq c_{0}+6$, so $y-x\geq 2$. This proves our claim that one of $q=q_{0}$ or $q=q_{0}-1$ will have $y-x\geq 2$.
Now taking either $q=q_{0}$ or $q=q_{0}-1$ so that $y-x\geq 2$, we have
$x=\left\lceil\frac{b+1}{q}\right\rceil\geq\frac{b}{q_{0}}\geq\sqrt{3.5}\sqrt{b}>\sqrt{3.4}\sqrt{b}\geq c,$
so $x\geq c+1$ because $x$ is an integer. Because $\sqrt{b}\geq 270$, we have
$q\geq q_{0}-1\geq\frac{\sqrt{b}}{\sqrt{3.5}}-2\geq\frac{\sqrt{b}}{\sqrt{3.8}}.$
Finally, we have
$\left\lceil\frac{x-1}{y-x}\right\rceil x$ $\leq\left(\frac{b}{2q}+1\right)\left(\frac{b}{q}+1\right)\leq\left(\frac{\sqrt{3.8}b}{2\sqrt{b}}+1\right)\left(\frac{\sqrt{3.8}b}{\sqrt{b}}+1\right)$
$\leq(0.975\sqrt{b}+1)(1.95\sqrt{b}+1)\leq(0.979\sqrt{b})(1.954\sqrt{b})\leq 1.92b\leq 2b+c+1.$
This completes the proof. ∎
Now we can prove the remaining Part 3 of Theorem 3.1.
Proposition 3.18.
If $c\geq 2$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
This follows directly from Proposition 3.13, Proposition 3.15, Lemma 3.14, and Lemma 3.17. ∎
We also prove that spiders with four legs are not $e$-positive.
Proof of Theorem 1.6. .
Because $S$ has four legs, we have $c\geq c_{1}+1\geq 2$. If $b\leq\frac{c^{2}}{2}$, the result follows from Parts 1, 2, and 3 of Theorem 3.1. If $b\geq\frac{c^{2}}{2}$, then
$n\geq 2b+c+1\geq c^{2}+c+1,$
and Zheng showed [spiderepos, Corollary 4.6] that such a spider $S$ is not $e$-positive. ∎
In order to prove Conjecture 1.1, it now remains to consider trees with a vertex of degree $4$ that is adjacent to a leaf, such as the spider $S(6,4,1,1)$. It turns out that $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$, which shows that Wolfgang’s necessary condition will not be enough. We conclude by showing that there are infinitely many such examples.
Figure 3. The spider $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$
$S(6,4,1,1)$
Proposition 3.19.
Let $m\geq 1$ be an integer. The spider $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 12m+1$.
Proof.
We claim that if $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ has a rearrangement $\alpha$ with $6m-1,6m\notin\text{set}(\alpha)$, then $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$. Label the vertices on the leg of length $26m-$ 2 as $u_{1},\ldots,u_{6m-2}$, moving in from the leaf, and label the vertices on the leg of length $6m$ as $v_{1},\ldots,v_{6m}$ similarly. Suppose that $\alpha$ is a rearrangement of $\lambda$ that satisfies
$\alpha_{1}+\cdots+\alpha_{i}\leq 6m-2\text{ and }\alpha_{1}+\cdots+\alpha_{i+1}\geq 6m+1.$
Then we can partition the vertices $\{u_{1},u_{2},\ldots,u_{\alpha_{1}+\cdots+\alpha_{i}}\}$ into paths of sizes $\alpha_{1},\ldots,\alpha_{i}$, we can partition the vertices $\{v_{1},v_{2},\ldots,v_{\alpha_{i+2}+\cdots+\alpha_{\ell}}\}$ into paths of sizes $\alpha_{i+2},\ldots,\alpha_{\ell}$, and the remaining vertices form a connected set of size $\alpha_{i+1}$. This establishes our claim.
Note that if $\lambda$ has at least two $1$’s or if $\lambda=(2,2,\ldots,2,1)$, then $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$, so we may assume that $\lambda$ has at most one $1$ and has a part $x\geq 3$. By our claim, it suffices to find a rearrangement $\alpha$ of $\lambda$ with $6m-1,6m\notin\text{set}(\alpha)$. Fix a rearrangement $\beta$ of $\lambda$. We may assume that $16m-$ 1 or $6m\in\text{set}(\beta)$, otherwise we could take $\alpha=\beta$. Letting $\beta^{*}=(\beta_{\ell},\ldots,\beta_{1})$ denote the reverse composition, we may assume that $16m-$ 1 or $6m\in\text{set}(\beta^{*})$, otherwise we could take $\alpha=\beta^{*}$. Now there are only a few cases to consider.
If $6m\in\text{set}(\beta)$ and $6m\in\text{set}(\beta^{*})$, we must have $\beta_{1}+\cdots+\beta_{i}=6m$, $\beta_{i+1}=1$, and $\beta_{i+2}+\cdots+\beta_{\ell}=6m$ for some $i$. By reversing, we may assume without loss of generality that some part $\beta_{k}=x\geq 3$ for $1\leq k\leq i$. Then the composition
$\alpha=(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{i},1,x,\beta_{i+2},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-x,6m-x+1,6m+1,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$.
If $6m-1\in\text{set}(\beta)$, $6m\notin\text{set}(\beta)$, and $6m\in\text{set}(\beta^{*})$, we must have $\beta_{1}+\cdots+\beta_{i}=6m-1$, $\beta_{i+1}=2$, and $\beta_{i+2}+\cdots+\beta_{\ell}=6m$ for some $i$. If there is some $\beta_{k}=x\geq 3$ for $1\leq k\leq i$, then the composition
$\alpha=(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{i},2,x,\beta_{i+2},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-x-1,6m-x+1,6m+1,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. Otherwise, the parts $(\beta_{1},\ldots,\beta_{i})$ consist of a single $1$ and $(3m-1)$ $2$’s, some $\beta_{k}=x\geq 3$ for $i+2\leq k\leq\ell$, and the composition
$\alpha=(2,2,\ldots,2,2,x,1,2,\beta_{i+2},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-2,6m-2+x,6m-1+x,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. If $6m\in\text{set}(\beta)$, $6m-1\in\text{set}(\beta^{*})$, and $6m\notin\text{set}(\beta^{*})$, we can reverse $\beta$ to reduce to the previous case.
Finally, if $6m-1\in\text{set}(\beta)$, $6m\notin\text{set}(\beta)$, $6m-1\in\text{set}(\beta^{*})$, and $6m\notin\text{set}(\beta^{*})$, then we must have $\beta_{1}+\cdots+\beta_{i}=6m-1$, $\beta_{i+1}=3$, and $\beta_{i+2}+\cdots+\beta_{\ell}=6m-1$ for some $i$. If $\beta_{k}=1$ for some $1\leq k\leq i$, then the composition
$\alpha=(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{i},3,1,\beta_{i+2},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-2,6m+1,6m+2,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. If $\beta_{k}=y\geq 4$ for some $1\leq k\leq i$, then the composition
$\alpha=(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{i},3,y,\beta_{i+2},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-1-y,6m+2-y,6m+2,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. Now we may assume that every $1\leq k\leq i$ has $\beta_{k}=2$ or $3$, and by reversing, that every $i+2\leq k\leq\ell$ also has $\beta_{k}=2$ or $3$. Because $16m-$ 1 is not divisible by either $2$ or $3$, there must be some $1\leq k\leq i$ with $\beta_{k}=3$ and some $i+2\leq k^{\prime}\leq\ell$ with $\beta_{k^{\prime}}=2$. Now the composition
$\alpha=(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{i},2,3,3,\beta_{i+2},\ldots,\beta_{k^{\prime}-1},\beta_{k^{\prime}+1},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-4,6m-2,6m+1,6m+4,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. This completes the proof. ∎
4. Acknowledgments
The author would like to think José Aliste-Prieto, Richard Stanley, Stephanie van Willigenburg, and Kai Zheng for helpful discussions.
[…]